# trans-subtitle [https://orcid.org/0000-0003-0875-8345] [https://orcid.org/0000-0002-6003-3168] [https://orcid.org/0009-0000-1789-4117] [https://orcid.org/0009-0000-8135-7260] [https://orcid.org/0000-0003-4632-5743] [https://orcid.org/0009-0000-5184-2695] [https://orcid.org/0000-0003-1324-1412]

### v1

@@ 2.3 NV Center Charge Properties
Surface defects, impurities, and subsurface traps induced by etching may also degrade the NV center charge state properties, 16, 42 resulting in lower charge initialization or poor spin contrast. We measure the steady state charge population of NV centers in the nanobeams and in the host diamond by recording the photon number histogram under weak ($\approx$3 µW) 594 nm orange illumination after conventional green initialization (Figure 3a). A double Poissonian distribution fit reveals the two charge states: NV- (higher photon counts) and NV0 (lower photon counts). We interrogate 116 NV centers in nanobeams of various widths and 148 NV centers in the host diamond and plot the resulting NV- population statistics in Figure 3b. The observed charge state distributions are very similar, with 50% and 45% of NV centers showing >60% NV- population in the nanobeams and the host diamond host, respectively.
@@
It has been reported that NV spin contrast can be compromised due to increased photoionization rates in the presence of imperfect surface termination even when the steady state NV- population is normal. 16 Therefore, we measure the time-dependent photoluminescence under orange illumination of variable power and fit the individual photoluminescence decays to an exponential function to extract the corresponding ionization rates. The rates vary quadratically with power as $aP^{2}$, where $a$ is the ionization coefficient and $P$ is the power, as expected for two-photon photoionization and recombination (see Figure S.5). The statistics of the ionization coefficients $a$ for 61 NV centers in the nanobeams of various widths and for 85 NV centers in the host diamond are plotted in Figure 3c. We find the mean ionization rate coefficients to be 5.4 Hz/µW2 and 3.6 Hz/µW2 for nanobeams and for the host diamond, respectively, with $\approx$20% of the NV centers in nanobeams displaying $a$ > 10 Hz/µW2. This faster ionization can be attributed to the increased optical excitation efficiency in the nanobeams (discussed in more detail in Section 2.4).
@@
To quantify the impact of this increased ionization rate on sensing performance, we measure the spin contrast by performing Rabi oscillations in a small magnetic field (25 G). The spin contrast is then calculated as $C=\frac{S_{0}-S_{1}}{S_{0}}$ where $S_{0}$ ($S_{1}$) is the signal measured after initializing the ground state spin of the NV center into a $m_{s}=0$ ($m_{s}=-1$) spin sublevel. The spin contrast distributions for 53 NV centers in the nanobeams of various widths and for 65 NV centers in the host diamond are shown in Figure 3d. 52% and 56% of NV centers in nanobeams and in the host diamond, respectively, are characterized by good spin contrast (>20%). We observe more NV centers with no spin contrast in the nanobeams (around 50% more), which we attribute to higher prevalence of NV centers with lower NV- population in this subset on the nanobeams (see Figure S.6). We conclude that the quality of the fabricated surfaces is compatible with adequate charge properties for nanoscale sensing.
@@ 2.4 Device Integration and Photonic Optimization
It has been shown that pulled glass rods can be utilized for pick and place transfer of diamond cantilevers. 43 To transfer our membranes, we use two glass micromanipulator probes, fabricated via micropipette pulling and mounted on motorized stages in a customized 2D-material transfer microscope (HQ Graphene). Using these probes, we first break the membrane tethers, then pick up the membranes, which are electrostatically held to the pipette tip. (Figure 4a). Using both manipulators, membranes are moved, flipped, and placed onto the surface of a sapphire substrate (Figure 4b). By applying mechanical pressure with a micromanipulator we can ensure that membranes make good contact with the substrate. For a clean sapphire surface, transferred membranes have sufficient contact to remain in place after wet processing with refluxing nitric, perchloric and sulfuric acid, as well as with HF, piranha solution (2:1 sulfuric acid and hydrogen peroxide), solvents such as acetone, IPA, and water, as well as blow drying under nitrogen.
@@ 3 Conclusions
We establish nanostructured diamond membranes that host state-of-the-art shallow NV centers as a practical platform for quantum sensing experiments. The membrane fabrication allows for long NV center spin coherence and stable charge dynamics, while structuring the membrane into nanobeams enables enhanced photon collection. We have demonstrated a device geometry that is compatible with polymer-free, pick-and-place mechanical transfer, which can be readily applied to sensing in a variety of superconducting, magnetic, and other condensed matter systems. In particular, this sensing geometry aids in the interrogation of optically opaque materials, which is a particularly challenging task for single-crystal diamond sensors. Additionally, enhanced collection efficiency and reduction in required excitation power improves measurement throughput and reduces optical illumination of the target material. Future potential extensions of these capabilities include patterned implantation of NV centers using a nanopatterned mask 44 or nanoimplantation techniques, 45 which would provide more reliable collection efficiency enhancement through control of the NV center lateral position, particularly with preferential dipole orientation alignment in (111)-oriented substrates. 46 This procedure could be extended to the fabrication of low-dimensional structures such as diamond nanosheets and nanowires, opening up a novel avenue for studying carrier47 and heat48 transport in diamond.
[…]
*Supporting Information Supporting Information is available from the Wiley Online Library or from the author
[…]
*Conflict of Interest
The authors declare no conflict of interest.
[…]
*Acknowledgments This work was primarily supported by the Gordon and Betty Moore Foundation (grant GBMF12237, DOI 10.37807) and by the National Science Foundation under Grant No. OMA-2326767. Low-damage diamond etching was primarily supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, under Award Number DE-SC0024537.
@@
We thank Alex Abulnaga for help with the initial FDTD code we used in this work. The authors acknowledge the use of Princeton’s Imaging and Analysis Center (IAC), which is partially supported by the Princeton Center for Complex Materials (PCCM), a National Science Foundation (NSF) Materials Research Science and Engineering Center (MRSEC; DMR-2011750) and John J. Schreiber for assistance in FIB and imaging. The devices in this work were fabricated at the Princeton Materials Institute Micro/Nanofabrication Center and Quantum Device Nanofabrication Laboratory - Princeton University. We also thank HQ graphene for their help and support building the micromanipulator transfer microscope.
[…]
*Author contributions A.C.P-T. developed the design and fabrication process, and performed measurements. A.L. carried out the spin and charge measurements. S.K. assisted with fabrication development and with T.B.M. conducted the modeling, photonic simulations, and PL measurements. M.L. performed the FIB cross-sectional analysis and assisted with fabrication. K.-H.C. and M.L. built the confocal microscope used for PL measurements. N.P.dL. supervised the project. A.C.P-T, A.L, S.K and N.P.dL wrote the manuscript.
[…]
*Financial Disclosure
None reported.
[…]
*Conflicts of Interest
The authors declare no conflicts of interest.
[…]
*Data Availability
@@
The data that support the findings of this study are available from the corresponding author upon reasonable request
[…]
*Keywords quantum sensing | diamond | nanofabrication | nitrogen vacancy | photonics | quantum sensor integration| membrane | nanobeam

### v2

@@ 2.3 NV Center Charge Properties
Surface defects, impurities, and subsurface traps induced by etching may also degrade the NV center charge state properties, [yuan2020charge, dhomkar2018charge] resulting in lower charge initialization or poor spin contrast. We measure the steady state charge population of NV centers in the nanobeams and in the host diamond by recording the photon number histogram under weak ($\approx$3 µW) 594 nm orange illumination after conventional green initialization (Figure 3a). A double Poissonian distribution fit reveals the two charge states: NV- (higher photon counts) and NV0 (lower photon counts). We interrogate 116 NV centers in nanobeams of various widths and 148 NV centers in the host diamond and plot the resulting NV- population statistics in Figure 3b. The observed charge state distributions are very similar, with 50% and 45% of NV centers showing ¿60% NV- population in the nanobeams and the host diamond host, respectively.
@@
It has been reported that NV spin contrast can be compromised due to increased photoionization rates in the presence of imperfect surface termination even when the steady state NV- population is normal. [yuan2020charge] Therefore, we measure the time-dependent photoluminescence under orange illumination of variable power and fit the individual photoluminescence decays to an exponential function to extract the corresponding ionization rates. The rates vary quadratically with power as $aP^{2}$, where $a$ is the ionization coefficient and $P$ is the power, as expected for two-photon photoionization and recombination (see Figure S.5). The statistics of the ionization coefficients $a$ for 61 NV centers in the nanobeams of various widths and for 85 NV centers in the host diamond are plotted in Figure 3c. We find the mean ionization rate coefficients to be 5.4 Hz/µW2 and 3.6 Hz/µW2 for nanobeams and for the host diamond, respectively, with $\approx$20% of the NV centers in nanobeams displaying $a$ ¿ 10 Hz/µW2. This faster ionization can be attributed to the increased optical excitation efficiency in the nanobeams (discussed in more detail in Section 2.4).
@@
To quantify the impact of this increased ionization rate on sensing performance, we measure the spin contrast by performing Rabi oscillations in a small magnetic field (25 G). The spin contrast is then calculated as $C=\frac{S_{0}-S_{1}}{S_{0}}$ where $S_{0}$ ($S_{1}$) is the signal measured after initializing the ground state spin of the NV center into a $m_{s}=0$ ($m_{s}=-1$) spin sublevel. The spin contrast distributions for 53 NV centers in the nanobeams of various widths and for 65 NV centers in the host diamond are shown in Figure 3d. 52% and 56% of NV centers in nanobeams and in the host diamond, respectively, are characterized by good spin contrast (>>20%). We observe more NV centers with no spin contrast in the nanobeams (around 50% more), which we attribute to higher prevalence of NV centers with lower NV- population in this subset on the nanobeams (see Figure S.6). We conclude that the quality of the fabricated surfaces is compatible with adequate charge properties for nanoscale sensing.
@@ 2.4 Device Integration and Photonic Optimization
It has been shown that pulled glass rods can be utilized for pick and place transfer of diamond cantilevers. [Maletinsky_tran] To transfer our membranes, we use two glass micromanipulator probes, fabricated via micropipette pulling and mounted on motorized stages in a customized 2D-material transfer microscope (HQ Graphene). Using these probes, we first break the membrane tethers, then pick up the membranes, which are electrostatically held to the pipette tip. (Figure 4a). Using both manipulators, membranes are moved, flipped, and placed onto the surface of a sapphire substrate (Figure 4b). By applying mechanical pressure with a micromanipulator we can ensure that membranes make good contact with the substrate. For a clean sapphire surface, transferred membranes have sufficient contact to remain in place after wet processing with refluxing nitric, perchloric and sulfuric acid, as well as with HF, piranha solution (2:1 sulfuric acid and hydrogen peroxide), solvents such as acetone, IPA, and water, as well as blow drying under nitrogen.
@@ 3 Conclusions
We establish nanostructured diamond membranes that host state-of-the-art shallow NV centers as a practical platform for quantum sensing experiments. The membrane fabrication allows for long NV center spin coherence and stable charge dynamics, while structuring the membrane into nanobeams enables enhanced photon collection. We have demonstrated a device geometry that is compatible with polymer-free, pick-and-place mechanical transfer, which can be readily applied to sensing in a variety of superconducting, magnetic, and other condensed matter systems. In particular, this sensing geometry aids in the interrogation of optically opaque materials, which is a particularly challenging task for single-crystal diamond sensors. Additionally, enhanced collection efficiency and reduction in required excitation power improves measurement throughput and reduces optical illumination of the target material. Future potential extensions of these capabilities include patterned implantation of NV centers using a nanopatterned mask [implant_aligned] or nanoimplantation techniques, [FIB_implant] which would provide more reliable collection efficiency enhancement through control of the NV center lateral position, particularly with preferential dipole orientation alignment in (111)-oriented substrates. [osterkamp_engineering_2019] This procedure could be extended to the fabrication of low-dimensional structures such as diamond nanosheets and nanowires, opening up a novel avenue for studying carrier[lozovoi2021optical] and heat[li2012thermal] transport in diamond.
Acknowledgements.
@@
This work was primarily supported by the Gordon and Betty Moore Foundation (grant GBMF12237, DOI 10.37807) and by the National Science Foundation under Grant No. OMA-2326767. Low-damage diamond etching was primarily supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, under Award Number DE-SC0024537. We thank Alex Abulnaga for help with the initial FDTD code we used in this work. The authors acknowledge the use of Princeton’s Imaging and Analysis Center (IAC), which is partially supported by the Princeton Center for Complex Materials (PCCM), a National Science Foundation (NSF) Materials Research Science and Engineering Center (MRSEC; DMR-2011750) and John J. Schreiber for assistance in FIB and imaging. The devices in this work were fabricated at the Princeton Materials Institute Micro/Nanofabrication Center and Quantum Device Nanofabrication Laboratory - Princeton University. We also thank HQ graphene for their help and support building the micromanipulator transfer microscope. A.C.P-T. developed the design and fabrication process, and performed measurements. A.L. carried out the spin and charge measurements. S.K. assisted with fabrication development and with T.B.M. conducted the modeling, photonic simulations, and PL measurements. M.L. performed the FIB cross-sectional analysis and assisted with fabrication. K.-H.C. and M.L. built the confocal microscope used for PL measurements. N.P.dL. supervised the project. A.C.P-T, A.L, S.K and N.P.dL wrote the manuscript.
@@
Data availability: The data that support the findings of this study are available from the corresponding author upon reasonable request.